Please help me solve this problem: Solve the linear equation: 18-(4m+6)=4+3m-13
Answer by fractalier(6550)

Okay, from
18-(4m+6)=4+3m-13
we can distribute the minus sign on the left and combine on the right, giving
18 - 4m - 6 = 3m - 9
12 - 4m = 3m - 9
Now add 4m and add 9 to both sides and get
21 = 7m
Now divide by 7
m = 3